a+b Cos x If tan () Q.9 tan tan tan then find the value of 0.
Answer #1

Clearly, tant (754- tan-(x) – tan-hly) | 1 + xy Therefore, tan Hisz) = tans (7771) = tan-12) - tap-111) ... (1) Likewise, 11 + tan- 1 + ... + ton- tan-11 1 I1 +1.2 1 +2.3 (1+ n(n+1)] ton- ( 1413 ) + tan (4:23) +-. *ter (itinima) = [ta6-4(2)- tan